OpenStudy (anonymous):

Solve for x: (x – 4)2 = 2

Parth (parthkohli):

Find the square root of both sides. \(\Large \color{Black}{\Rightarrow x - 4 = \sqrt{2} }\) Add 4 to both sides. \(\Large \color{Black}{\Rightarrow x = \sqrt{2} + 4 }\)